OpenStudy (anonymous):

The volume of a sphere is 4,000π m3. What is the surface area of the sphere to the nearest square meter? 50,265 m2 181 m2 1,307 m2 2,614 m2

OpenStudy (anonymous):

Start by finding the radius of the sphere. Plug into the formula \[\huge \frac{4}{3} \pi r^3=V\]

OpenStudy (wolf1728):

r^3 = V•(3/4) / PI